# Survey on the Evaluation of Generative Models in Music

### v1

@@ Abstract.
Research on generative systems in music has seen considerable attention and growth in recent years. A variety of attempts have been made to systematically evaluate such systems. We provide an interdisciplinary review of the common evaluation targets, methodologies, and metrics for the evaluation of both system output and model usability, covering subjective and objective approaches, qualitative and quantitative approaches, as well as empirical and computational methods. We discuss the advantages and challenges of such approaches from a musicological, an engineering, and an HCI perspective.
@@ 1. Introduction
In recent years, advances in system architecture and training methodologies of generative systems in machine learning have led to increasingly powerful systems that have been applied to a variety of tasks and use cases. Particularly prominent are systems in natural language processing (e.g., (Brown et al., 2020)) and image models (e.g., (Karras et al., 2019; Ramesh et al., 2021)). Generative systems for music, while not featured as prominently in the media, have also seen considerable progress. While the concept of computer-generated music goes back to the mid-20th century (Hiller and Isaacson, 1959; Supper, 2001) and has been actively researched since then (Cope, 2000; Fernandez and Vico, 2013), there has been a rapid increase in interest and publications in the past five years (Civit et al., 2022), mostly driven by advances in neural approaches. Inspired by Herremans et al. (2017), we categorize generative systems for music according to (i) architecture, (ii) output, and (iii) input or control. [^1] References will be limited to a select number of representative systems. A meta-review by Civit et al. (2022) lists the prevalent architectures with decreasing number of occurrences as Recurrent Neural Networks (RNNs) (Hadjeres et al., 2017; Mao et al., 2018), Feed-Forward Networks (FF) (Yang et al., 2017; Dong et al., 2018), Variational Audio Encoders (VAEs) (Roberts et al., 2018; Dhariwal et al., 2020), evolutionary algorithms (Stoltz and Aravind, 2019; De Prisco et al., 2020), and Transformer-based approaches (Huang et al., 2019; Donahue et al., 2019; Copet et al., 2023), sometimes combined with Generative Adversarial Networks (GANs) (Yang et al., 2017; Dong et al., 2018). Details of these typical neural architectures for music generation can be found in Briot et al. (2020). A key distinction of output types of generative systems in music is whether the output is audio (Dieleman et al., 2018; Copet et al., 2023) or symbolic (e.g., MIDI, MusicXML, etc.) (Huang et al., 2019; Roberts et al., 2018). The length of the output can also vary: it might range from as short as a single note or event in the case of audio synthesizers (Engel et al., 2017, 2020; Kong et al., 2020) over phrases (Naruse et al., 2022; Huang and Yang, 2020) to complete musical pieces (Dhariwal et al., 2020; Lam et al., 2023). Furthermore, systems might generate a single-voiced melody (Yang et al., 2017; Hakimi et al., 2020) or a polyphonic output with multiple voices (Hadjeres et al., 2017; Herremans and Chew, 2019; Oore et al., 2020; Huang and Yang, 2020). The input of a system depends very much on the design goals. A system might require no input at all (Pasini and Schlüter, 2022), a few parameters for conditioning (Chen et al., 2020; Eisenbeiser, 2020), a text prompt (Copet et al., 2023; Agostinelli et al., 2023), a melody to be harmonized (Hadjeres et al., 2017; Yeh et al., 2021), or a musical phrase to be continued (Huang et al., 2019; Huang and Yang, 2020) or inpainted (Pati et al., 2019; Marafioti et al., 2021; Araneda-Hernandez et al., 2023). While this variety of approaches and the multitude of available studies imply rapid progress, this progress is hard to quantify, and there is evidence that the quality improvement might not as dramatic as the number of publications suggests (Yin et al., 2023).
@@
These challenges have led to a large variety in approaches to system evaluation with a multitude of different evaluation targets, methodologies, and metrics. Inter-study inconsistencies in evaluation make a comparison of research results essentially impossible. If results cannot be compared, do not sufficiently reflect the actual quality of a system, or have been acquired in very different settings, the notion of progress in this field becomes questionable, as we cannot measure progress without relevant, commonly used metrics. Despite these problems being recognized as important challenges (Yang and Lerch, 2020; Briot et al., 2020; Liang, 2023), no general solutions have been proposed, and evaluation still seems to be largely neglected or treated as an afterthought. For instance, Civit et al. (2022) provide a meta-review of generative music systems but only mention evaluation in passing. While Bandi et al. (2023) present a dedicated evaluation section in their extensive review of generative systems, music is unfortunately not discussed. One laudable exception is Ji et al. (2023) who summarize some objective and subjective approaches to evaluation.
Therefore, the goal of this article is to provide an accessible, interdisciplinary overview on current empirical and quantitative approaches to the evaluation of generative systems in music, providing an in-depth discussion of evaluation targets and methodologies for the assessment of the output of generative systems as well as the user interaction with such methods (rather than other assessment targets such as sociological implications, etc.). In order to do so, we first introduce a comprehensive overview of the dimensions or targets to be evaluated in Sect. 2, followed by a description of methodologies and metrics to evaluation of system output and user interaction in Sects. 3 and 4, respectively. We conclude the paper with a discussion on challenges and future work in Sect. 5 and final remarks in Sect. 6.
@@ 2. Evaluation targets
The main goals of evaluating a machine learning system are (i) to determine if or to what degree a system works as intended and (ii) to (quantitatively) compare it with other systems. These goals, however, can have many facets as an evaluation can focus on different targets. Wang et al. (2024a) group evaluation targets into “data-quality evaluation” and “property-controllability evaluation,” the latter focusing on specific output properties that are implicitly or explicitly controlled. Pasquier et al. (2017) lists the following aspects of a generative system to be assessed: quality, creativity, believability, complexity, robustness, and reliability. In this work, we propose to group the main evaluation targets into system output and model use while acknowledging that there are aspects of the model itself and the process of creation that could be subject of evaluation as well.
@@
As music is a fundamental form of human creativity (Kozbelt, 2019), it is also critical to evaluate model use — how easy to use, enjoyable, and engaging models are for people when they make music whether they be hobbyists, music students, or professional musicians. A model may produce high quality output, but if it is unusable then it has little value for making music. Evaluating the use of models draws on research in Human-Computer Interaction (HCI) (Preece et al., 2015) to assess whether a model is, e.g., usable by people (Nielsen, 1994), or whether it provokes feelings of surprise (Caramiaux and Fdili Alaoui, 2022) or reflection (Ford and Bryan-Kinns, 2023).
@@ 2.1.1. Aesthetic and musical qualities
Despite the fact that the assessment of aesthetic quality is possibly the most commonly stated evaluation target in the literature, it arguably is the most challenging to operationally define. While aesthetic quality (as applied to a musical artifact) is a catch-all phrase encompassing many attributes such as balance, complexity, novelty, etc. (Brattico and Jacobsen, 2009), it is most commonly evaluated only in the singular dimension of subjective preference, i.e., how much a listener “likes” a piece of music, and/or how “interesting” it is (Dervakos et al., 2021). However, individual differences are known to impact such aesthetic judgments as well, in particular, those that arise from the acquisition of musical expertise (Pearce, 2015).
@@
In order to compare across studies, it is important that researchers explicitly operationally define the traits they aim to measure, not only for themselves, but to potential participants as well. For instance, Brattico and Jacobsen (2009) mention the important distinction between affective responses (those that induce or modulate emotions or mood), hedonic responses (those that modulate reward; likes and dislikes), and aesthetic responses, which typically refer to inherent style-relevant attributes which lend the artifact beauty, elegance, or coherence. To complicate matters, in addition to style-relevant traits, the medium of creation (e.g., score versus audio), the cultural context (e.g., what is valued inside versus outside the culture), and the caliber or quality of the object or its execution are all important criteria that factor into the equation of evaluating aesthetic goals (Galanter, 2012). For instance, music generation that outputs a score or transcription may be evaluated on its adherence to various compositional norms, such as adherence to an appropriate vocabulary and grammar, the arrangement and organization of musical ideas, and the use of variation and repetition, to name a few (Sturm and Ben-Tal, 2017). On the other hand, the output rendered as an audio recording may be evaluated based on the execution of the performance in relation to parameters such as the authenticity or “humanness” of the performance, the expressivity or dynamicism, potentially in addition to factors related to the underlying composition itself. These criteria are extremely multifaceted. For instance, the analysis and assessment of music performance is a research field in itself (Katayose et al., 2012; Lerch et al., 2019, 2020).
@@
While the assessment of aesthetics typically involves human evaluation, there have been attempts at computational aesthetic evaluation. Galanter (2012) gives an overview of such methods through 2012, noting that “computational aesthetic evaluation is an extremely difficult problem,” and that it frequently “leads to deep philosophical waters regarding phenomenology and consciousness.” Models that have been proposed and used in a musical context, have typically relied on evaluation by adherence to some set of statistical measures or proportions. Some AI models that generate an artistic work may be designed so as to implicitly include the goals of such fitness metrics, leading to increased diversity, for example. As pointed out by Galanter (2012) and others, however, “[c]reating evolutionary diversity and dynamics via artificial aesthetics foreign to our human sensibility is one thing. Appealing to human aesthetics is quite another.” Other computational avenues for evaluation have included models based complexity — frequently drawing on Berlyne’s theory of aesthetics, or Shannon information theory (Galanter, 2012).
@@ 2.1.3. Originality
The originality of the output is a common concern when evaluating a generative system. We interpret originality in three ways. First, originality is a concern with respect to plagiarism. Commonly, a generative system is expected to create novel output that does not replicate the training data. Second, the diversity of the model output should match the diversity of the training data in all relevant dimensions. This validates the training success. Last but not least, we may want to measure the creativity of the system output.
@@
Modern machine learning systems require a potentially massive amount of training data. In most music scenarios, the (partial) reproduction of memorized training data might be regulated by laws. Although systems are generally trained with the intent to generate novel artifacts, they might memorize individual training samples and reproduce them during inference (van den Burg and Williams, 2021; Golda et al., 2024). When that happens, claims of plagiarism can arise similar to when a human composer copies musical ideas from existing works. Examples for memorized output from generative systems in other fields include reproduced names and email addresses for language models (Carlini et al., 2021; Huang et al., 2022), and reproduced images for image generation models (Carlini et al., 2023). The analysis of plagiarism in music is more complicated, so cases are not often directly or easily identifiable. Therefore, most of the current discussion of plagiarism in music focuses on the use of copyrighted data for training and whether this establishes an infringement of copyright (Sunray, 2021). Currently, we see the first lawsuits around this topic reaching the courts (Kinsella, 2024). As such potential infringements and associated questions on liability are not necessarily easily addressed and answered, assessing the probability of a generative system reproducing memorized content can be part of an overall evaluation strategy.
@@
Creativity is a notoriously difficult to define concept. As pointed out by Jordanous (2012), there are not even standard definitions of creativity within US or UK law, “despite the need to detect the presence of creativity for legal reasons.” Even more stubborn are the issues that arise in attempting to assess creativity, as pointed out by Rohrmeier (2022): “It is hard to assess something that one cannot define, and this reflects down to the difficulties in evaluating the success of models of general creativity without resorting to the ’oracle’ of human evaluators.” While Rohrmeier is referring to the assessment of models of general creativity, the same issues arise in the assessment of creative artifacts or model outputs, regardless of the methods of creation (Carnovalini and Rodà, 2020). Despite the fact that creativity is recognized as an ill-defined phenomenon, studies continue to attempt to evaluate it, as it remains a crucial component of artistic creation (Kvak, 2022). However, in many (if not most) modes of musical creation, one is bound by a set of rules or constraints (Chu et al., 2022). Again, Rohrmeier (2022) raises the question of creativity in the context of style imitation where such rules and constraints can, in some cases, be extreme: “What do we mean by ’creativity’, and how do we relate novelty, innovation or transformation with the concept? For instance, are models ’creative’ that generate jazz lead sheets, chorales in Bach’s style, Indian tabla, or Balinese Gamelan? Is style replication ’creative’?” As pointed out by Agres et al. (2016), creativity can be evaluated in a limited sense in a highly constrained context —such as the harmonization of a melody in a strict chorale style— as the comparison between human and computational ability to solve a set of problems. This general logic can be extended to theoretically any aspect of creativity. However, even in only evaluating the artifact of a style-imitation task, this procedural or problem-solving definition becomes a slippery slope towards pure determinism, which has been argued to be the opposite of (or, at least, hindering to) creativity (Rohrmeier, 2022; Zimmermann, 1996). In addition, many attempts to define creativity (e.g., (Boden, 1998; Jordanous, 2012; Cope, 2005; Ritchie, 2007)) commonly include the notions of value (either aesthetic or utilitarian), the combination or connection of ideas or phenomena, and exploration and transformation within some conceptual space, all of which are missed in the constrained problem-solving definition of style imitation described above.

### v2

@@ Abstract.
Research on generative systems in music has seen considerable attention and growth in recent years. A variety of attempts have been made to systematically evaluate such systems. We present an interdisciplinary review of the common evaluation targets, methodologies, and metrics for the evaluation of both system output and model use, covering subjective and objective approaches, qualitative and quantitative approaches, as well as empirical and computational methods. We examine the benefits and limitations of these approaches from a musicological, an engineering, and an HCI perspective.
@@ 1. Introduction
In recent years, advances in system architecture and training methodologies of generative systems in machine learning have led to increasingly powerful systems that have been applied to a variety of tasks and use cases. Particularly prominent are systems in natural language processing (e.g., (Brown et al., 2020)) and image models (e.g., (Karras et al., 2019; Ramesh et al., 2021)). Generative systems for music, while not featured as prominently in the media, have also seen considerable progress. Although the concept of computer-generated music dates back to the mid-20th century (Hiller and Isaacson, 1959; Supper, 2001) and has been actively researched since then (Cope, 2000; Fernandez and Vico, 2013), there has been a rapid increase in interest and publications in the past five years (Civit et al., 2022), mostly driven by advances in neural approaches. Inspired by Herremans et al. (2017), we categorize generative music systems based on (i) architecture, (ii) output, and (iii) input or control. [^1] References will be limited to a select number of representative systems. A meta-review by Civit et al. (2022) lists the prevalent architectures with decreasing number of occurrences as Recurrent Neural Networks (RNNs) (Hadjeres et al., 2017; Mao et al., 2018), Feed-Forward Networks (FF) (Yang et al., 2017; Dong et al., 2018), Variational Audio Encoders (VAEs) (Roberts et al., 2018; Dhariwal et al., 2020), evolutionary algorithms (Stoltz and Aravind, 2019; De Prisco et al., 2020), and Transformer-based approaches (Huang et al., 2019; Donahue et al., 2019; Copet et al., 2023), sometimes combined with Generative Adversarial Networks (GANs) (Yang et al., 2017; Dong et al., 2018), with Diffusion networks (Evans et al., 2024, 2025; Liu et al., 2023; Chen et al., 2024; Schneider et al., 2024; Ning et al., 2025), or with Rectified Flows (Li et al., 2025). Details of many typical neural architectures for music generation can be found in Briot et al. (2020).
A key distinction of output types of generative systems in music is whether the output is audio (Dieleman et al., 2018; Copet et al., 2023; Agostinelli et al., 2023) or symbolic (e.g., MIDI, MusicXML, etc.) (Huang et al., 2019; Roberts et al., 2018). The length of the output can also vary: it might range from as short as a single note or event in the case of audio synthesizers (Engel et al., 2017, 2020; Kong et al., 2020b) over phrases (Naruse et al., 2022; Huang and Yang, 2020) to complete musical pieces (Dhariwal et al., 2020; Lam et al., 2023). Furthermore, systems might generate a single-voiced melody (Yang et al., 2017; Hakimi et al., 2020) or a polyphonic output with multiple voices (Hadjeres et al., 2017; Herremans and Chew, 2019; Oore et al., 2020; Huang and Yang, 2020).
The input of a system depends very much on the design goals. A system might require no input at all (Pasini and Schlüter, 2022), a few parameters for conditioning (Chen et al., 2020; Eisenbeiser, 2020), a text prompt (Copet et al., 2023; Agostinelli et al., 2023), a melody to be harmonized (Hadjeres et al., 2017; Yeh et al., 2021), or a musical phrase to be continued (Huang et al., 2019; Huang and Yang, 2020) or inpainted (Pati et al., 2019; Marafioti et al., 2021; Araneda-Hernandez et al., 2023). While this variety of approaches and the multitude of available studies imply rapid progress, this progress is hard to quantify, and there is evidence that the quality improvements might not be as dramatic as the number of publications suggests (Yin et al., 2023).
@@
These challenges have led to a large variety in approaches to system evaluation with a multitude of different evaluation targets, methodologies, and metrics. Inter-study inconsistencies in evaluation make the comparison of research results essentially impossible. If results cannot be compared, do not sufficiently reflect the actual quality of a system, or have been acquired in very different settings, the notion of progress in this field becomes questionable, as we cannot measure progress without relevant, commonly used metrics. Despite these problems being recognized as important challenges (Yang and Lerch, 2020; Briot et al., 2020; Liang, 2023), no general solutions have been proposed, and evaluation still seems to be largely neglected or treated as an afterthought. For instance, Civit et al. (2022) provide a meta-review of generative music systems but only mention evaluation in passing. Zhao et al. (2025) review prompt-based generative music systems but refer only to the evaluation of creativity as an unsolved challenge. While Bandi et al. (2023) present a dedicated evaluation section in their extensive review of generative systems, music is unfortunately not discussed. The only two exceptions are Ji et al. (2023) and Wang et al. (2024c), summarizing some objective and subjective approaches to evaluation.
Therefore, the goal of this article is to provide an accessible, interdisciplinary overview on current empirical and quantitative approaches to the evaluation of generative systems in music. Figure 1 gives a summary and accessible flow chart of the approaches presented. The article provides an in-depth discussion of evaluation targets and methodologies for the assessment of the output of generative systems as well as the user interaction with such methods (rather than other assessment targets such as sociological implications, etc.). In order to do so, we first introduce a comprehensive overview of the dimensions or targets to be evaluated in Sect. 2, followed by a description of methodologies and metrics to evaluation of system output and user interaction in Sects. 3 and 4, respectively. We conclude with a discussion on challenges and future directions in Sect. 5 and final remarks in Sect. 6.
Figure 1. Structure of the presented evaluation approaches and methodologies.
[…]
Structural flow-chart overview of the topics covered in the paper, starting with evaluation targets (musical qualities, audio fidelity, originality, and conditioning for output evaluation and ease of use, creative experience, user impression, and musical practice for evaluation of usability and user experience. Subsequent steps outline the detailed methodologies usable for assessment.
@@ 2. Evaluation targets
The main goals of evaluating a machine learning system are (i) determining whether the system works as intended and to what degree and (ii) comparing it (quantitatively) with other systems. These goals, however, can have many facets as an evaluation can focus on different targets. Wang et al. (2024a) group evaluation targets into “data-quality evaluation” and “property-controllability evaluation,” the latter focusing on specific output properties that are implicitly or explicitly controlled. Pasquier et al. (2017) lists the following aspects of a generative system to be assessed: quality, creativity, believability, complexity, robustness, and reliability. In this work, we propose to group the main evaluation targets into system output and model use while acknowledging that there are aspects of the model itself and the process of creation that could be subject of evaluation as well.
@@
Given that music is a fundamental form of human creativity (Kozbelt, 2019), it is also critical to evaluate model use — how easy to use, enjoyable, and engaging models are for people when they make music, whether they be hobbyists, music students, or professional musicians. A model may produce high quality output, but if it is unusable then it has little value for making music. Evaluating the use of models draws on research in Human-Computer Interaction (HCI) (Preece et al., 2015) to assess whether a model is, e.g., usable by people (Nielsen, 1994), or provokes surprise (Caramiaux and Fdili Alaoui, 2022) or reflection (Ford and Bryan-Kinns, 2023).
@@ 2.1.1. Aesthetic and musical qualities
Despite the fact that the assessment of aesthetic quality is possibly the most commonly stated evaluation target in the literature, it arguably is the most challenging to operationally define. While aesthetic quality (as applied to a musical artifact) is a catch-all phrase encompassing many attributes such as balance, complexity, novelty, etc. (Brattico and Jacobsen, 2009), it is most commonly evaluated only in the singular dimension of subjective preference, i.e., how much a listener “likes” a piece of music, and/or how “interesting” it is (Dervakos et al., 2021). However, individual differences, particularly those that arise from the acquisition of musical expertise, are known to impact such aesthetic judgments (Pearce, 2015).
@@
In order to compare across studies, it is important that researchers explicitly operationally define the traits they aim to measure, not only for themselves, but to potential participants as well. For instance, Brattico and Jacobsen (2009) mention the important distinction between affective responses (those that induce or modulate emotions or mood), hedonic responses (those that modulate reward; likes and dislikes), and aesthetic responses, which typically refer to inherent style-relevant attributes which lend the artifact beauty, elegance, or coherence. To complicate matters, in addition to style-relevant traits, the medium of creation (e.g., score versus audio), the cultural context (e.g., what is valued inside versus outside the culture), and the caliber or quality of the object or its execution are all important criteria that factor into the equation of evaluating aesthetic goals (Galanter, 2012). For instance, music generation that outputs a score or transcription may be evaluated on its adherence to various compositional norms, such as adherence to an appropriate vocabulary and grammar, the arrangement and organization of musical ideas, and the use of variation and repetition, to name a few (Sturm and Ben-Tal, 2017). On the other hand, the output rendered as an audio recording may be evaluated based on the execution of the performance in relation to parameters such as the authenticity or “humanness” of the performance, the expressivity or dynamicism, potentially in addition to factors related to the underlying composition itself. These criteria are highly multifaceted and context-dependent. For instance, the analysis and assessment of music performance is a research field in itself (Katayose et al., 2012; Lerch et al., 2019, 2020).
@@
While the assessment of aesthetics typically involves human evaluation, there have been attempts at computational aesthetic evaluation. Galanter (2012) gives an overview of such methods through 2012, noting that “computational aesthetic evaluation is an extremely difficult problem,” and that it frequently “leads to deep philosophical waters regarding phenomenology and consciousness.” Models that have been proposed and used in a musical context, have typically relied on evaluation by adherence to some set of statistical measures or proportions. Some AI models that generate an artistic work may be designed so as to implicitly include the goals of such fitness metrics, leading to increased diversity, for example. As pointed out by Galanter (2012), however, “[c]reating evolutionary diversity and dynamics via artificial aesthetics foreign to our human sensibility is one thing. Appealing to human aesthetics is quite another.” Other computational avenues for evaluation have included models based complexity — frequently drawing on Berlyne’s theory of aesthetics, or Shannon information theory (Galanter, 2012).
@@ 2.1.3. Originality
The originality of the output is a common concern when evaluating a generative system. We interpret originality in three ways. First, originality is a concern with respect to plagiarism. Commonly, a generative system is expected to create novel output that does not replicate the training data. Second, the diversity of the model output should match the diversity of the training data in all relevant dimensions. This validates the effectiveness of the training. Last but not least, we may want to measure the creativity of the system output.
@@
Modern machine learning systems require a potentially massive amount of training data. In most music scenarios, the (partial) reproduction of memorized training data might be regulated by laws. Although systems are generally trained with the intent to generate novel artifacts, they might memorize individual training samples and reproduce them during inference (van den Burg and Williams, 2021; Golda et al., 2024). When that happens, claims of plagiarism can arise similar to when a human composer copies musical ideas from existing works. Examples for memorized output from generative systems in other fields include reproduced names and email addresses for language models (Carlini et al., 2021; Huang et al., 2022), and reproduced images for image generation models (Carlini et al., 2023). The analysis of plagiarism in music is more complicated, so cases are not often directly or easily identifiable. Therefore, most of the current discussion in music focuses on copyright (Lee, 2024) and the use of copyrighted data for training and whether this establishes an infringement of copyright (Sunray, 2021; Desai and Riedl, 2024). Currently, we see the first lawsuits around this topic reaching the courts (Kinsella, 2024). As such potential infringements and associated questions on liability are not necessarily easily addressed and answered, assessing the probability of a generative system reproducing memorized content can be part of an overall evaluation strategy.
@@
Creativity is a notoriously difficult to define concept. As pointed out by Jordanous (2012), there are not even standard definitions of creativity within US or UK law, “despite the need to detect the presence of creativity for legal reasons.” Even more stubborn are the issues that arise in attempting to assess creativity, as pointed out by Rohrmeier (2022): “It is hard to assess something that one cannot define, and this reflects down to the difficulties in evaluating the success of models of general creativity without resorting to the oracle’ of human evaluators.” While Rohrmeier is referring to the assessment of models of general creativity, the same issues arise in the assessment of creative artifacts or model outputs, regardless of the methods of creation (Carnovalini and Rodà, 2020). Despite the fact that creativity is recognized as an ill-defined phenomenon, studies continue to attempt to evaluate it, as it remains a crucial component of artistic creation (Kvak, 2022). However, in many (if not most) modes of musical creation, one is bound by a set of rules or constraints (Chu et al., 2022). Again, Rohrmeier (2022) raises the question of creativity in the context of style imitation where such rules and constraints can, in some cases, be extreme: “What do we mean by creativity’, and how do we relate novelty, innovation or transformation with the concept? For instance, are models ’creative’ that generate jazz lead sheets, chorales in Bach’s style, Indian tabla, or Balinese Gamelan? Is style replication creative’?” As pointed out by Agres et al. (2016), creativity can be evaluated in a limited sense in a highly constrained context —such as the harmonization of a melody in a strict chorale style— as the comparison between human and computational ability to solve a set of problems. This general logic can be extended to theoretically any aspect of creativity. However, even in only evaluating the artifact of a style-imitation task, this procedural or problem-solving definition becomes a slippery slope towards pure determinism, which has been argued to be the opposite of (or, at least, hindering to) creativity (Rohrmeier, 2022; Zimmermann, 1996). In addition, many attempts to define creativity (e.g., (Boden, 1998; Jordanous, 2012; Cope, 2005; Ritchie, 2007)) commonly include the notions of value (either aesthetic or utilitarian), the combination or connection of ideas or phenomena, and exploration and transformation within some conceptual space, all of which are missed in the constrained problem-solving definition of style imitation described above.